OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 2x - cos x

zepdrix (zepdrix):

Recall the `Sine Double Angle Identity`:\[\large \color{green}{\sin 2x=2\sin x \cos x}\] Understand how you can use this to deal with the sin2x in your problem? :)

OpenStudy (anonymous):

here are the options 2 sin x cos2x sin x cos x (2 sin x - 1) 2 sin x

OpenStudy (anonymous):

no i don't understand?

zepdrix (zepdrix):

\[\large \color{green}{\sin 2x}-\cos x\]Using the identity I pasted above we can rewrite our equation as,\[\large \color{green}{2\sin x\cos x}-\cos x\] From here we can factor out a cosx from each term. What would that give you? :o

OpenStudy (anonymous):

no the 3rd one

OpenStudy (anonymous):

right?

zepdrix (zepdrix):

lol i have a feeling you're just guessing XD

OpenStudy (anonymous):

No i really think it's the 3rd one

zepdrix (zepdrix):

Ok, good job :D ya it's the third one.